Session: May 1999 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (G.Maccaferri at Sun Jun 13 12:05:43 METDST 1999) Success

Noto (C.Trigilio at Mon Jun 14 08:20:41 METDST 1999) Success

Effelsberg (A. Kraus at Mon Jun 14 09:16:30 METDST 1999) Minor failures
Problems with the positioning of tape head one have been reported between UT 0230 - 0238 and UT 0752 - 0800.

Onsala (A. Polatidis at Mon Jun 14 16:45:09 METDST 1999) Success

Torun (K. Borkowski at Tue Jun 15 15:44:29 METDST 1999) Minor failures
Around 7 UTC it was discovered that there is a constant offset of 0.14 deg in azimuth introduced by a calibration procedure before the experiment started. So until this time there was a considerable loss of signal.

Jodrell (Alastair Gunn at Fri Jun 18 11:00:10 METDST 1999) Success
No known problems.

Cambridge (Alastair Gunn at Fri Jun 18 11:00:38 METDST 1999) Minor failures
Problem communicating with MERLIN LOs at start of experiment, followed by FS crash, means data bad until 03:40 UT scan.

Westerbork (Tony Foley at Fri Jun 18 16:22:34 METDST 1999) Minor failures
Used 13 telescopes in array. Parity checks procedure during the observation gave high rates- Socorro please check.

Calibration (Jean\-Francois Desmurs at Mon Jun 21 16:14:53 METDST 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may99 directory under the name: es030b.antab. Read the comments in the file to get further information. (Rem : Cm no Tsys values.)